With the estimates shown below, Sarah needs to determine the trade-in (replacement) value of machine X that will render its AW equal to that of machine Y at an interest rate of 9% per year. Determine the replacement value.

Machine X

Machine Y

Market Value, $

?

92,000

Annual Cost, $ per Year

−55,500

−40,000 for year 1,increasing by 2000 per year thereafter.

Salvage Value

15,500

17,000

Life, Years

3

5

The replacement value is $ ________.
Answer #1

ANSWER:

We will equate the aw of machine x and machine y equal.

i = 9%

aw of machine x = market value(a/p,i,n) + annual cost + salvage value(a/f,i,n)

aw of machine x = market value(a/p,9%,3) - 55,500 + 15,500(a/f,9%,3)

aw of machine x = market value * 0.3951 - 55,500 + 15,500 * 0.3051

aw of machine x = 0.3951 market value - 55,500 + 4,729.05

aw of machine x = 0.3951 market value - 50,771

pw of machine y = market value + annual cost(p/a,i,n) + increase in cost(p/g,i,n) + salvage value(p/f,i,n)

pw of machine y = -92,000 - 40,000(p/a,9%,5) - 2,000(p/g,9%,5) + 17,000(p/f,9%,5)

pw of machine y = -92,000 - 40,000 * 3.89 - 2,000 * 7.111 + 17,000 * 0.6499

pw of machine y = -92,000 - 155,600 - 14,222 + 9,098.6

pw of machine y = -252,723

aw of machine y = pw of machine y(a/p,i,n)

aw of machine y = -252,723(a/p,9%,5)

aw of machine y = -252,723 * 0.2571

aw of machine y = -64,975.2

aw of machine x = aw of machine y

0.3951 market value - 50,771 = -64,975.2

0.3951 market value = -64,975.2 + 50,771

0.3951 market value = -14,204.2

market value = -14,204.2 / 0.3951

market value = -35,950.9

so the market value is $35,950.9 that will make the aw of both the machines equal.